OpenStudy (kise_ryouta):

Solve for x. x+3.2=−3.5

Vocaloid (vocaloid):

subtract 3.2 from both sides

Vocaloid (vocaloid):

x + 3.2 - 3.2 = ?

OpenStudy (kise_ryouta):

-6.7

Vocaloid (vocaloid):

nope, try again

Vocaloid (vocaloid):

what is 3.2 - 3.2?

OpenStudy (kise_ryouta):

-6.4

Vocaloid (vocaloid):

nope

OpenStudy (kise_ryouta):

0 ?

Vocaloid (vocaloid):

good, 0 so what does x + 3.2 - 3.2 equal?

OpenStudy (kise_ryouta):

-6.4

Vocaloid (vocaloid):

no, we already said that 3.2 - 3.2 was 0 so what is x + 3.2 - 3.2?

OpenStudy (kise_ryouta):

0

Vocaloid (vocaloid):

what about the x?

OpenStudy (kise_ryouta):

it stays?

Vocaloid (vocaloid):

good, so what does x + 3.2 - 3.2 equal?

OpenStudy (kise_ryouta):

x=0

Vocaloid (vocaloid):

x does not equal 0

Vocaloid (vocaloid):

think about it: we start with x add 3.2 take away 3.2 what's left?

OpenStudy (kise_ryouta):

-3.5 is left

Vocaloid (vocaloid):

where are you getting 3.5????

OpenStudy (kise_ryouta):

Solve for x. x+3.2=−3.5 <-------

Vocaloid (vocaloid):

no, we're just talking about x + 3.2 - 3.2

OpenStudy (kise_ryouta):

x is left right

Vocaloid (vocaloid):

good!

Vocaloid (vocaloid):

x + 3.2 - 3.2 = x next step!

Vocaloid (vocaloid):

since we subtracted 3.2 from the left side, we must also subtract it from the right side what is -3.5 - 3.2?

OpenStudy (kise_ryouta):

-6.7?

Vocaloid (vocaloid):

good job! therefore, x = -6.7

OpenStudy (kise_ryouta):

OHHH thank Ow< can you help me with more pls